OpenStudy (anonymous):

some what simple parametric equations question find the points on the cruve where the tanget is horizontal or vertical. x=t^3-3t y=t^2-3 i know how to find the derivative, i get (2t)/(3t^2-3) the tangent should be vertical when t= negative or postive 1 right? but my book says negative or postive 2 for some reason. and i get that the tanget is horzontal at x=0, since that would make the numerator 0.

OpenStudy (anonymous):

sec, i just realized i mistyped

OpenStudy (anonymous):

it should be horizontal at the point x=0, and vertical at x= plus or minus 2.

OpenStudy (anonymous):

You know the equation for \(dy/dx\) right? So find out when numerator is 0 (horizontal) And when denominator is 0 (vertical)

OpenStudy (anonymous):

When both are 0 then.. well things get tricky.

OpenStudy (anonymous):

denominator =0 when x =negative of postive 1. but in my book it says negative or positive 2.

OpenStudy (anonymous):

negative or positive*

OpenStudy (anonymous):

did you write the problem correctly?

OpenStudy (anonymous):

x=t^3-3t and y=t^2-3 is exactly how its written in the book

OpenStudy (anonymous):

maybe i took the dy/dx incorrectly, but i dont see any mistakes. its just a matter of taking the derivative of each parametric equation, then putting the dy over dx. (dy/dx)

OpenStudy (anonymous):

maybe the book is wrong haha? doubt it though ._.

OpenStudy (anonymous):

Hmmmm, what happens at 2 and -2 anyway?

OpenStudy (anonymous):

not entirely sure. but according to the book thats where the tangent line is vertical.

OpenStudy (anonymous):

Hmmm, maybe you are answering in terms of \(t\) but should answer in terms of \(x\)

OpenStudy (anonymous):

What is \( x(t=-1)\) and \(x(t=1)\)

OpenStudy (anonymous):

when t=1, x=-2 and when t=-1, x =2

OpenStudy (anonymous):

ic now, guess i had to plug values back in to parametric equations

OpenStudy (anonymous):

ty